Jack the jaguar can run for only 1.0 min at 13.0 m/s before he has to take a break. Zeke the zebra can gallop at 7.25 m/s for 5.0 min. How long does Zeke have to run to elude Jack? Can Jack catch Zeke for lunch if they are 350 m apart?
1
Expert's answer
2018-09-10T12:59:08-0400
1)
d_J=v_J t_J=(13)(60)=780 m.
The time needed is
t=d_J/v_Z =780/7.25 s=780/7.25 1/60 min=1.8 s
2)
d_J-D=780-350=430 m.
Zeke the zebra can travel at this minute:
d_Z=7.25(60)=435 m.
d_Z>d_J-D
Thus, Jack cannot catch Zeke for lunch if they are 350 m apart.